A cavity contains black body radiation in equilibrium at temperature T. The specific heat per unit volume of the photon gas in the cavity is of the form Cv=γT3,where γ is a constant. The cavity is expanded to twice its original volume and then allowed to equilibrate at the same temperature T. The new internal energy per unit volume is
1
Expert's answer
2020-10-01T09:58:39-0400

"c_v=\gamma T^{3}"


also


"C_V=(\frac{du}{dT})_u" ,therefore,


"\gamma T^{3}=(\frac{du}{dT})"


"\gamma T^3 dT=du"


Integrating both sides


"\int\gamma T^3 dT=\int du"


"\gamma \frac{T^4}{4}=u" (Internal energy per unit volume)